7. Find average kinetic energy of a molecule in a room at 25.0 degrees C? (assume air is an ideal gas)
Answer #1

Solution ) According to Kinetic Molecular Theory, the average kinetic energy of gas molecules is a function only of temperature.

The formula is

KE=3/2 kT

where T is the Kelvin temperature and k is Boltzmann's constant.

KE=3/2×1.381×10⁻²³J⋅K⁻¹×298.15K = 6.176 × 10⁻²¹ J

For a mole of molecules, the average kinetic energy is

((6.176×10⁻²¹J)/1molecule)×((6.022×10²³molecules)/1mol)

= 3.719 ×10³ J/mol = 3.719  kJ/mol

Hello dear if any other issues with this solutions please informing me... for my request at least before your negative response please comment I'll help you ..we look forward for your happiness ....thank you ..keep smiling .. :)